OpenStudy (anonymous):

What is the value of f[g(6)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (anonymous):

to find f(g(x)), you need to plug in g(x) everywhere where f(x) has x. f(g(x)) = 2(2x - 5) + 1 = 4x + 11 f(g(6)) = 34

OpenStudy (anonymous):

love you